### v3

@@ 4.1 Analysis of Anthropometric Measurements
Eleven anthropometric measures were used to compute descriptive statistics using Microsoft Excel 2021. The following values were estimated to create a standard database, as shown in Table 4: minimum (min), maximum (max), mean, standard deviation, and various percentile values (5th, 50th, and 95th percentiles). The anthropometric measurements of the male and female participants in this study were computed independently.
Table 4: Descriptive statistics of the anthropometric measurements taken for both male and female participants (in mm)
Measurements Gender Min Max Mean Standard Deviation Percentile
5th 50th 95th
PH Male 414 489 444.37 13.37 422.65 444.96 466.49
female 373 447 415.1 13.5 395.2 414.6 438.4
SEH Male 188 308 234.7 16.73 205.99 235.05 261.23
female 191 282 231.3 16.7 200.6 230.3 257.6
BPL Male 402 498 454.03 18.17 422.91 454.1 483.72
female 394 478 447.2 17.3 412.8 448.2 471
AL Male 331 394 364.1 10.77 346.35 363.86 382.29
female 322 366 342.4 9.6 327 342.1 355.9
TT Male 131 178 156.14 7.76 143.46 155.99 168.68
female 126 175 145.1 9.7 129.3 144.6 161.4
SSH Male 470 556 511.78 14.9 488.6 512.03 539.42
female 447 522 488 13.3 467 488.9 509.2
SCH Male 484 540 511.7 10.22 495.19 511.42 528.01
female 469 524 493.6 10.8 475.1 493.8 510.1
EFL Male 426 471 450.46 7.67 437.08 450.21 462.3
female 385 428 406.9 9.5 391.7 407.5 421.4
SEB Male 396 492 447.47 15.36 424.46 447.56 470.46
female 390 464 422.9 17.2 396.1 421.2 452.9
HB Male 328 382 350.53 9.4 335.31 350.23 365.34
female 344 390 366.2 8.5 353.3 367.2 379.4
BKL Male 496 548 521.47 9.54 505 522.23 536.1
female 475 546 509 13.7 489.2 509 529.9
The normal bell-shaped curves observed in Figures 4(a) and 4(b) indicate the reliability and consistency of the collected data. In statistics, a normal distribution is often considered a hallmark of reliable data, as it suggests that the measurements are evenly distributed around the mean, with fewer extreme values and more values clustered around the center. The presence of these curves in the data for both sexes indicates that the measurements were taken accurately and consistently without significant outliers or inconsistencies. This reliability is crucial in ergonomic research, as it ensures that the collected data accurately represent the population and can be used to make informed decisions about furniture design and ergonomics. Additionally, the normal distributions in both male and female datasets suggest that the collected measurements represent the general population rather than being skewed or biased toward one gender. This further enhances the reliability and validity of the dataset and supports its applicability in various settings and industries.
(a) Distribution of male anthropometric data
(b) Distribution of female anthropometric data
The Spearman correlation method was used to analyze the correlation between all 11 anthropometric measurements. As depicted in Figure 6, the heatmap generated from the analysis represents positive correlations in red and negative correlations in blue. Notably, the elbow fingertip length exhibited a medium positive correlation with PH (0.44), AL (0.46), and SCH (0.43). Conversely, HB was moderately negatively correlated with EFL (-0.43). SEH and BPL showed negligible correlation with other anthropometric measurements, with values close to zero. The Spearman correlation heatmap values ranged from -1 to +1, indicating the strength and direction of the correlations between the variables.
Figure 6: Correlation between all 11 anthropometric measurements (for both male and female combined data).
@@ 4.2 Analysis of Existing Furniture Dimensions
Table 5 lists the 11 existing furniture dimensions of the computer furniture that were measured. Type-1 indicates a table with a non-adjustable chair, whereas type-2 indicates a table with an adjustable chair.
@@ 4.3 Mismatch Analysis
To evaluate the degree of mismatch between the current library furniture and body measurements, equations 2 through 12 were applied. A mismatch was found if the equation was not satisfied for equations with one-way relationships. Three classes were established to quantify the match or mismatch for equations with two-way relations. A match was found when the anthropometric measurement was in between the ranges. When the relation’s highest limit is less than the anthropometric measurement, a low mismatch is shown. A high mismatch is observed when the relation’s lowest limit is greater than the measurement. To find a mismatch, the writers looked into 11 aspects of computer furniture. Individual matches or mismatches were identified by comparing the dimensions of the furnishings with the body measurements of each participant. Finally, a percentage representing the measured match or mismatch for the sample population was given.
In Table 6, for the first type of furniture (non-adjustable chair with non-adjustable table), the mismatch level is notably higher for male body measurements than females in parameters, such as SH is 74.17%, SW is 59.92%, BW is 60.33% and SD is 5.75% higher whereas BH is 54.75%, and STH is 6.17% lower. Only one parameter, UEB, was found to be a 100% match for both genders. The results were particularly poor for the table-related parameters, indicating that furniture was poorly fitted for its users in these dimensions. All the parameters are compared in Figure 7 to better understand the scenario.
Table 6: Summary of the level of mismatch and match percentages of existing furniture dimensions of 1st type of furniture (non-adjustable chair and table)
@@
Figure 7: Mismatch percentages of existing furniture dimensions of 1st type of furniture (non-adjustable chair and table).
The results for the second type of furniture (adjustable chair and non-adjustable table) in Table 7 show significantly lower mismatch levels than the first type of furniture for SH, SW, BW, STC, and UTH. For this particular furniture set, BW, UEB, and STH showed zero mismatches and needed no improvements. However, the outcomes for the table-related parameters remain subpar, indicating that furniture is still not well-suited for its users in these dimensions. Figure 8 compares all the parameters more explicitly.
@@
Figure 8: Mismatch of existing furniture dimensions of 2nd type of furniture (adjustable chair and non-adjustable table).
4.4 ANOVA Test
@@
A one-way ANOVA test was used to determine whether there were any statistically significant differences between the means of three or more independent groups. In this study, furniture dimensions were tested against anthropometric dimensions to determine if there were any significant differences. The level of significance ($\alpha$) value was set at 5%, meaning that a result was considered statistically significant if the p-value was $\leq 5\%$. The result was not statistically significant if the p-value was $>5\%$.
The ANOVA test resulted in a critical value (F-value) and a probable value (p-value). The F-value measures the variance between groups compared with the variance within the groups. The p-value is the probability of observing a result as extreme as that obtained, assuming that the null hypothesis is true. If the p-value was less than the $\alpha$ value (5% in this case), the null hypothesis was rejected, and it was concluded that there was a statistically significant difference between the groups. If the p-value was greater than the $\alpha$ value, the null hypothesis was not rejected, and it was concluded that there was no statistically significant difference between the groups.
@@
In this study, ANOVA was used to determine whether there were any statistically significant differences in furniture dimensions and anthropometric measurements, both for males and females. The results showed that SH, SW, BH, UEB, STH, STC, and TL of furniture dimensions were rejected for non-adjustable chairs and tables, meaning there were statistically significant differences between the groups. In contrast, the SD and STH were accepted, indicating that there were statistically no significant differences between the groups (Table 8). For the adjustable chair non-adjustable table, we considered both the highest and lowest limits of SH, STH, and STC. If either the highest or lowest limit is accepted for males and females, we consider the parameter accepted. According to the ANOVA test for adjustable chairs with non-adjustable tables, only SH, SD, and STH dimensions were accepted. The rest of the dimensions, SW, BH, BW, UEB, STC, and TL were rejected (Table 9).
@@
  • $H_{0}$(null hypothesis): There is no effect or difference between the anthropometric measurements and existing furniture dimensions.
@@
  • $H_{1}$(alternative hypothesis): There is an effect or difference between the anthropometric measurements and existing furniture dimensions.
Table 8: ANOVA test results for non-adjustable furniture dimensions
5th 50th 95th
SH vs PH Male Observed 422.65 444.96 466.49 0.71 0.447 Accept
Expected 443.79 457.2 470.61
Female Observed 395.17 414.58 438.4 7.83 0.049 Reject
Expected 443.79 457.2 470.61
SW vs HB Male Observed 335.31 350.23 365.34 11.19 0.029 Reject
Expected 376.98 393.7 410.42
Female Observed 353.28 367.24 379.45 4.51 0.101 Accept
Expected 375.97 393.7 411.43
SD vs BPL Male Observed 422.91 454.1 483.72 5.38 0.081 Accept
Expected 388.67 406.4 424.13
Female Observed 412.78 448.24 470.99 3.61 0.13 Accept
Expected 388.67 406.4 424.13
BH vs SSH Male Observed 488.6 512.03 539.42 173.83 0 Reject
Expected 294.62 304.8 314.98
Female Observed 467.05 488.85 509.16 184.73 0 Reject
Expected 294.62 304.8 314.98
BW vs HB Male Observed 335.31 350.23 365.34 0.28 0.625 Accept
Expected 346.88 355.6 364.32
Female Observed 353.28 367.24 379.45 1.48 0.29 Accept
Expected 346.88 355.6 364.32
UEB vs SCH Male Observed 495.19 511.42 528.01 70.93 0.001 Reject
Expected 392.32 406.4 420.48
Female Observed 475.12 493.8 510.12 44.56 0.003 Reject
Expected 392.32 406.4 420.48
STH vs SEH Male Observed 205.99 235.05 261.23 0.17 0.694 Accept
Expected 230.81 241.3 251.79
Female Observed 200.65 230.26 257.61 0.45 0.538 Accept
Expected 230.81 241.3 251.79
STC vs TT Male Observed 143.46 155.99 168.68 58.13 0.002 Reject
Expected 80.32 88.9 97.48
Female Observed 129.33 144.58 161.45 28.58 0.006 Reject
Expected 80.32 88.9 97.48
TL vs BKL Male Observed 505 522.23 536.1 13.8 0.021 Reject
Expected 473.68 482.6 491.52
Female Observed 489.25 509.05 529.86 4.37 0.105 Accept
Expected 473.68 482.6 491.52
Table 9: ANOVA test results for adjustable chair with non-adjustable table dimensions
5th 50th 95th
SH (lowest limit) vs PH Male Observed 422.65 444.96 466.49 0.527 0.508 Accept
Expected 410.21 431.8 453.39
Female Observed 395.17 414.58 438.4 0.796 0.423 Accept
Expected 410.21 431.8 453.39
SH (highest limit) vs PH Male Observed 422.65 444.96 466.49 19.804 0.011 Reject
Expected 506.73 533.4 560.07
Female Observed 395.17 414.58 438.4 35.008 0.004 Reject
Expected 506.73 533.4 560.07
SW vs HB Male Observed 335.31 350.23 365.34 45.836 0.002  Reject
Expected 434.34 457.2 480.06
Female Observed 353.28 367.24 379.45 35.436 0.004 Reject
Expected 434.34 457.2 480.06
SD vs BPL Male Observed 422.91 454.1 483.72 1.023 0.369 Accept
Expected 410.21 431.8 453.39
Female Observed 412.78 448.24 470.99 0.337 0.593 Accept
Expected 410.21 431.8 453.39
BH vs SSH Male Observed 488.6 512.03 539.42 148.401 0 Reject
Expected 289.56 304.8 320.04
Female Observed 467.05 488.85 509.16 149.574 0 Reject
Expected 289.56 304.8 320.04
BW vs HB Male Observed 335.31 350.23 365.34 9.223 0.039 Reject
Expected 374.02 393.7 413.39
Female Observed 353.28 367.24 379.45 3.926 0.119 Accept
Expected 374.02 393.7 413.39
UEB vs SCH Male Observed 495.19 511.42 528.01 48.612 0.002 Reject
Expected 386.08 406.4 426.72
Female Observed 475.12 493.8 510.12 31.274 0.005 Reject
Expected 386.08 406.4 426.72
STH (lowest limit) vs SEH Male Observed 205.99 235.05 261.23 0.101 0.766 Accept
Expected 217.17 228.6 240.03
Female Observed 200.65 230.26 257.61 0.003 0.962 Accept
Expected 217.17 228.6 240.03
STH (highest limit) vs SEH Male Observed 205.99 235.05 261.23 26.745 0.007 Reject
Expected 313.69 330.2 346.71
Female Observed 200.65 230.26 257.61 28.057 0.006 Reject
Expected 313.69 330.2 346.71
STC (lowest limit) vs TT Male Observed 143.46 155.99 168.68 61.03 0.001 Reject
Expected 90.49 95.25 100.01
Female Observed 129.33 144.58 161.45 26.571 0.007 Reject
Expected 90.49 95.25 100.01
STC (highest limit) vs TT Male Observed 143.46 155.99 168.68 19.526 0.012  Reject
Expected 187.01 196.85 206.69
Female Observed 129.33 144.58 161.45 22.616 0.009 Reject
Expected 187.01 196.85 206.69
TL vs BKL Male Observed 505 522.23 536.1 16.011 0.016 Reject
Expected 434.34 457.2 480.06
Female Observed 489.25 509.05 529.86 8.739 0.042 Reject
Expected 434.34 457.2 480.06
@@ 4.5 Proposed Dimensions of Furniture
Autodesk Fusion 360 software was used for designing furniture. Table 10 outlines the suggested dimensions for furniture, categorized into chairs and tables, based on principles that consider the anthropometric measurements of the participants. The proposed UTH for both table types is 645mm as shown in Figure 9. The SH is suggested to be between 400 mm and 450 mm for adjustable chairs, based on the 5th percentile of female popliteal height with 30 mm shoe clearance (see Figure 9(a)). The recommended SH is 430 mm for non-adjustable chairs, as shown in Figure 9(b). The proposed SW is 425 mm, based on the 95th percentile of female HB. The recommended SD is 385 mm, determined by the 5th percentile of female BPL. BH is suggested to be 350 mm, considering the 5th percentile of female SSH. The recommended BW is 390 mm, based on the 95th percentile of male SEB. UEB is proposed to be 465 mm, considering the minimum subscapular height of female participants. The suggested UTH is 645 mm, considering the 5th percentile of male SEH.
Figure 9: Proposed UTH for both types of furniture in mm.
@@
Table 10: Proposed dimensions for both types of furniture sets
Dimensions For type-1 furniture set For type-2 furniture set
SH 430 400 - 450
SW 425 425
SD 385 385
BH 350 350
BW 390 390
UEB 465 465
STH 260 235 - 310
STC 200 95.25 - 200
UTH 645 645
For non-adjustable chairs and tables, Table 11 presents the mismatch percentages of the proposed furniture measurements. This is also shown graphically in Figure 11. The SH against PH showed 5% and 12.5% for males and females, respectively. The SW against HB showed zero mismatch for Male and only 1.25 lower mismatch for females. The SD against BPL showed 7.33% high mismatch and 0.33 % low mismatch for males, whereas 0% high mismatch and 1.25% low mismatch for females. It is necessary to reduce the high mismatch for females as much as possible while designing the chair’s SD. The BH, BW, UEB, and STC showed a 100% match for both males and females. It was not possible to design a single non-adjustable chair and table set for 100% match for both males and females. So, the dimensions were maintained to balance a reduced mismatch percentage for both males and females for STH and UTH dimensions. For adjustable chairs with tables, it was possible to match the 100% furniture dimensions for both males and females. SH and STH showed better results for both males and females due to the chair’s height adjustability. SW, SD, and UTH showed similar results to non-adjustable chairs with table furniture sets. Though adjustable chairs showed improved results in some scenarios, we considered both types of furniture to be financially feasible and avoided design complexity.
VDT should be placed according to recent studies; computer monitor placement should be based on the needs of the particular user and should be performed within a modest height range below eye level with a head tilt that is physiologically beneficial (Woo et al.,, 2016). OSHA (2008) suggests that monitors should be positioned 50–100 cm away from users while maintaining a viewing angle of $15^{\circ}$$20^{\circ}$ (Weidling and Jaschinski,, 2015).
Using a keyboard and mouse is a frequent task in a computer lab; it should be placed in the frequent or regular zone of the Barnes and Squires Work Envelope Model. This indicates that the keyboard and mouse should be placed within the 394 mm of table depth and 1194 mm of table length (Norman,, 2013).
Compared to the previous mismatch percentages in university furniture, the proposed dimensions exhibited reduced mismatches for both males and females. The suggested measurements are more compatible than the previous dimensions, leading to improved comfort levels and reduced risks of MSDs for users.
@@
Figure 11: Mismatch of existing furniture dimensions of 1st type of furniture (non-adjustable chair and non-adjustable table).
Table 12 presents the mismatch percentages of the proposed furniture measurements for adjustable chairs and tables. This is also shown in Figure 12. Compared to previous mismatch percentages in university furniture, the proposed dimensions exhibited reduced mismatches for both males and females. The suggested measurements are more compatible than the previous dimensions, leading to improved comfort levels and reduced risks of MSDs for users.
@@
Figure 12: Mismatch of existing furniture dimensions of 2nd type of furniture (adjustable chair and non-adjustable table).

### v4

@@ 4.1 Analysis of Anthropometric Measurements
Eleven anthropometric measures were used to compute the descriptive statistics. The following values were estimated to create a standard database (Table A1): minimum (min), maximum (max), mean, standard deviation, and various percentile values (5th, 50th, and 95th percentiles). Anthropometric measurements of the male and female participants in this study were computed independently.
4.2 Assessment of Intra- and Inter-Class Reliability
Intra-class reliability was assessed by comparing the measurements taken by the same measurer on two separate occasions, whereas inter-class reliability was assessed by comparing measurements taken by different measures. Intra- and inter-class reliability for the anthropometric measurements was assessed using a two-way mixed effects model, where people’s effects were considered random, and the measure effects were fixed. The assessment of intra- and inter-class reliability is crucial for ensuring the precision and consistency of our data.
In this study, we obtained high intra-class correlation coefficients ($ICC_{1}$) for both single measures ($ICC_{1}$ = 0.973, 95% CI: 0.966 - 0.979) and average measures ($ICC_{1}$ = 0.986, 95% CI: 0.983 - 0.989). These findings indicate a high level of agreement and consistency between the repeated measurements. The inter-class reliability ($ICC_{2}$) was assessed by comparing the measurements taken by different researchers. Similarly, we found high $ICC_{2}$ values for both single measures ($ICC_{2}$ = 0.944, 95% CI: 0.931-0.956) and average measures ($ICC_{2}$ = 0.971, 95% CI: 0.964 - 0.977). These results indicate a high level of agreement and consistency between measurements taken by different researchers.
@@ 4.3 Analysis of Existing Furniture Dimensions
Table 1 lists the 11 existing furniture dimensions of the measured computer furniture. Type-1 indicates a table with a non-adjustable chair, whereas type-2 indicates a table with an adjustable chair.
@@ 4.4 Mismatch Analysis
To evaluate the degree of mismatch between the current library furniture and body measurements, equations 2 through 12 were applied. To find a mismatch, the writers examined into 11 aspects of computer furniture. Individual matches or mismatches were identified by comparing the dimensions of the furnishings with the body measurements of each participant. Finally, a percentage representing the measured match or mismatch for the sample population was presented.
Table 2: Summary of the level of mismatch and match percentages of existing furniture dimensions of 1st type of furniture (non-adjustable chair and non-adjustable table)
@@
4.5 Statistical Test
@@
One-way ANOVA was used to determine whether there were any statistically significant differences between the means of three or more independent groups. In this study, the furniture dimensions were tested against anthropometric dimensions to determine any significant differences. The level of significance ($\alpha$) was set at 5%, meaning that the result was considered statistically significant if the p-value was $\leq 5\%$. The result was not statistically significant if the p-value was $>5\%$. The F-value measures the variance between groups compared with the variance within groups. The p-value is the probability of observing a result as extreme as obtained, assuming that the null hypothesis is true. If the p-value was less than the $\alpha$ value (5% in this case), the null hypothesis was rejected, and it was concluded that there was a statistically significant difference between the groups. If the p-value was greater than the $\alpha$ value, the null hypothesis was not rejected, and it was concluded that there was no statistically significant difference between the groups.
@@
The results showed that SH, SW, BH, UEB, STH, STC, and TL of furniture dimensions were rejected for the type-1 furniture set, meaning there were statistically significant differences between the groups. In contrast, SD and STH were accepted, indicating that there were no statistically significant differences between the groups (Table A2). For the adjustable chair non-adjustable table, we considered both the highest and lowest limits of SH, STH, and STC. If the highest or lowest limit is accepted for males and females, we consider the parameter accepted. Only the SH, SD, and STH dimensions were accepted according to the ANOVA test for adjustable chairs with non-adjustable tables. The remaining dimensions, SW, BH, BW, UEB, STC, and TL, were rejected (Table A3).
@@
  • $H_{0}$ (null hypothesis): There is no effect or difference between anthropometric measurements and existing furniture dimensions.
@@
  • $H_{1}$ (alternative hypothesis): There is an effect or difference between anthropometric measurements and existing furniture dimensions.
@@ 4.6 Proposed Dimensions of Furniture
Table 4 presents the proposed dimensions for the two types of furniture sets designed for university computer laboratories, along with a comparison with relevant International Standards (ISO/EN) for ergonomic furniture design. The dimensions include SH, SW, SD, BH, BW, UEB, STH, STC, and UTH. The type-1 and type-2 furniture sets have slightly different proposed dimensions, particularly for SH, STH, and STC. The table also indicates each dimension’s corresponding International Standards (ISO/EN), providing a reference for optimal ergonomic design in university computer laboratories. The SH is suggested to be between 400 mm and 450 mm for adjustable chairs( Figure 4(a)). The recommended SH is 430 mm for non-adjustable chairs, as shown in Figure 4(b). The proposed SW is 425 mm, SD is 385 mm, BH is suggested to be 350 mm, BW is 390 mm, UEB is proposed to be 465 mm, and UTH is 645 mm.
@@
Table 4: Proposed dimensions of furniture sets and comparison with International Standards
Dimensions Type-1 Proposed Dimensions (mm) Type-2 Proposed Dimensions (mm) Proposed dimensions according to International Standards (mm)
SH 430 400 - 450 410 - 490 (ISO 5970:1979)
SW 425 425 440 - 480 (EN 1729-1:2020)
SD 385 385 390 - 450 (EN 1729-2:2012)
BH 350 350 340 - 400 (EN 1729-2:2012)
BW 390 390 370 - 430 (ISO 5970:1979)
UEB 465 465 450 - 530 (EN 1335-1:2000)
STH 260 235 - 310 260 - 340 (EN 1729-1:2016-02)
STC 200 95.25 - 200 95 - 225 (EN 1729-1:2016-02)
UTH 645 645 625 - 705 (ISO 5970:1979)
For non-adjustable chairs and tables, Table 5 presents the mismatch percentages of the proposed furniture measurements. The SH against PH showed 5% and 12.5% for males and females, respectively. The SW against HB showed zero mismatch for males and only 1.25 lower mismatch for females. The SD against BPL showed 7.33% high mismatch and 0.33 % low mismatch for males, whereas 0% high mismatch and 1.25% low mismatch for females. It is necessary to reduce the high mismatch for females as much as possible when designing a chair’s SD. BH, BW, UEB, and STC showed a 100% match for both males and females.
Compared with the previous mismatch percentages in university furniture, the proposed dimensions exhibited reduced mismatches for both males and females. The suggested measurements were more compatible than the previous dimensions, leading to improved comfort levels and reduced risks of MSDs for users.
@@
Table 6 presents the mismatch percentages of the proposed furniture measurements for the adjustable chairs and tables. Compared to the previous mismatch percentages in university furniture, the proposed dimensions exhibited reduced mismatches for males and females. The suggested measurements were more compatible than the previous dimensions, leading to improved comfort levels and reduced risks of MSDs for users.
@@
5 Discussion
The normal bell-shaped curves observed in Figures 1(a) and 1(b) indicate the reliability and consistency of the collected 11 anthropometric data collected. In statistics, a normal distribution is often considered a hallmark of reliable data, as it suggests that the measurements are evenly distributed around the mean, with fewer extreme values and more values clustered around the center. The presence of these curves in the data for both sexes indicated that the measurements were accurate and consistent without significant outliers or inconsistencies. This reliability is crucial in ergonomic research, as it ensures that the collected data accurately represent the population and can be used to make informed decisions about furniture design and ergonomics. Additionally, the normal distributions in both male and female datasets suggest that the collected measurements represent the general population rather than being skewed or biased toward one gender. This further enhances the reliability and validity of the dataset and supports its applicability to various settings and industries.
The high intra- and inter-class reliabilities of our study indicate that our anthropometric measurements are highly reliable and consistent. These findings enhance the validity and reliability of our data and provide confidence in the accuracy of our measurements.
In Table 2, for the first type of furniture, the mismatch level is notably higher for male body measurements than for females in parameters, such as SH was 74.17%, SW was 59.92%, BW was 60.33%, and SD was 5.75% higher whereas BH was 54.75%, and STH was 6.17% lower. Only one parameter, UEB, was found to be a 100% match for both sexes. The results were particularly poor for table-related parameters, indicating that furniture was poorly fitted for users in these dimensions. The results for the second type of furniture in Table 3 show significantly lower mismatch levels than those for the first type of furniture for SH, SW, BW, STC, and UTH. BW, UEB, and STH showed zero mismatches for this furniture set, and no improvements were required. However, the mismatch percentage is still sufficiently large enough for the remaining eight parameters, which still require improvements. However, the outcomes for the table-related parameters remain subpar, indicating that furniture is still not well-suited for users in these dimensions.
After the statistical test, the decisions were more understandable, and 11 furniture dimensions were proposed based on 11 anthropometric data collected. The proposed dimensions for UTH showed a 100% match for both furniture sets. The SH was proposed based on the 5th percentile of the female popliteal height with 30 mm shoe clearance. The proposed SW is 425 mm, based on the 95th percentile of HB in females. The recommended SD is 385 mm, as determined by the 5th percentile of female BPL. BH is suggested to be 350 mm, considering the 5th percentile of female SSH. The recommended BW is 390 mm, based on the 95th percentile of male SEB. The UEB is proposed to be 465 mm, considering the minimum subscapular height of the female participants. The suggested UTH is 645 mm, considering the 5th percentile of male SEH.
Designing a non-adjustable chair with a non-adjustable table set with a 100% match for both males and females was impossible. Therefore, the dimensions were maintained to balance a reduced mismatch percentage for both males and females for the STH and UTH dimensions. For adjustable chairs with tables, it was possible to match the 100% furniture dimensions for males and females. SH and STH showed better results for males and females because of the chair’s height adjustability. SW, SD, and UTH showed results similar to those of non-adjustable chairs with table furniture sets. Although adjustable chairs showed improved results in some scenarios, we considered both types of furniture to be financially feasible and avoided design complexity.
The proposed dimensions for both types of furniture followed international standards except for SD and SW. The SD and SW are slightly less than those of the European Committee for Standardization (2016) [30], but these values are very close to the standard values. The reason behind this is that the proposed dimensions are based on our collected samples that contained more male data and fewer female data. The HB is generally lower in males than in females.
VDT should be placed according to recent studies; computer monitor placement should be based on the needs of a particular user and should be performed within a modest height range below eye level with a head tilt that is physiologically beneficial [54]. Occupational Safety and Health Administration (2008) [27] suggests that monitors should be positioned 50–100 cm away from users while maintaining a viewing angle of $15^{\circ}$$20^{\circ}$ [55].
Using a keyboard and mouse is a frequent task in a computer lab, and it should be placed in the frequent or regular zone of the Barnes and Squires Work Envelope Model. This indicates that the keyboard and mouse should be placed within the 394 mm of table depth of 1194 mm and a table length [56].
The research is centered on the design of two sets of furniture identified within computer laboratories across various departments at KUET. Although the dimensions may exhibit variability, both types of furniture share similarities and are commonly employed in computer laboratories across diverse university settings. The proposed dimensions for these furniture sets were tailored to accommodate the anthropometric measurements of their users without introducing design intricacies or cost escalation.
The practical benefits of this research are significant for educational institutions, furniture manufacturers, and students. Inconsistencies in furniture dimensions were found by comprehensively analyzing anthropometric measurements and their correlation with furniture dimensions, which offers valuable insights into designing more comfortable and ergonomically sound furniture for university students. For educational institutions, this study suggests that investing in furniture tailored to students’ body dimensions can contribute to a more conducive learning environment. Comfortable and ergonomic furniture can help students maintain better posture, reduce the risk of MSDs, and improve their overall well-being and academic performance. Furniture manufacturers can benefit from this research by using the proposed dimensions as a guideline for designing products that are better suited to the needs of students. By incorporating these dimensions into their designs, manufacturers can create furniture that is more comfortable, functional, and attractive for their target market.
Students stand to gain the most from this research as they are the end users of furniture. By using furniture designed with body dimensions in mind, students can enjoy a more comfortable and ergonomic learning experience. This can improve focus, concentration, and productivity and ultimately enhance academic performance.
This research also has managerial implications for educational institutions and stakeholders furnishing university spaces. Educational institutions can use these findings to inform strategic planning, prioritizing investments in ergonomic furniture tailored to students’ anthropometric measurements. By incorporating the research findings into procurement requirements, institutions can ensure that they purchase furniture that meets students’ ergonomic needs, thereby increasing their satisfaction and well-being. Based on this research, facility managers can optimize classroom and computer lab layouts, ensuring that furniture accommodates students’ anthropometric measurements. This promotes a more comfortable and efficient learning environment and enhances student satisfaction, well-being, and academic performance. Educating students about the importance of ergonomic furniture and proper posture can prevent MSDs and related issues and improve student satisfaction, well-being, and academic performance. Establishing guidelines for furniture based on anthropometric criteria ensures compliance with relevant regulations and standards and provides students with safe, comfortable, and ergonomic learning environments. Furthermore, investing in ergonomic furniture tailored to anthropometric measurements can enhance health, comfort, productivity, and sustainable practices, leading to cost savings, an improved reputation, and legal compliance.
The limitations of this research include the reliance on a specific population (university students); the population size was 380. Although the number is statistically justified from a practical point of view, it should be increased further in future research. The exclusion of certain anthropometric measurements and the lack of consideration of other factors influencing furniture design (e.g., cultural preferences and environmental constraints) is another limitation of our research. Additionally, the proposed furniture dimensions may not be universally applicable and require further validation in other settings or populations. Future research could explore the implementation of the proposed dimensions in real-world classroom settings and assess their long-term impacts on student comfort and well-being using a larger population size. In addition, measuring the human body repeatedly would show slightly different results every time, which was neglected in our research. Further studies should investigate the effects of different furniture designs on student performance and productivity in academic settings.